OpenStudy (anonymous):

Sketch the graph of the equation (x−2)2 + (y+3)2 = 36 .

OpenStudy (jdoe0001):

http://www.mathwarehouse.com/geometry/circle/images/equation-of-circle/general-formula-equation-of-circle.png \(\bf (x−2)^2 + (y+3)^2 = 36\implies (x−2)^2 + (y+3)^2 = 6^2\)